OpenStudy (anonymous):

@hoblos : If y varies directly as x and y = 14 when x = 8, find x when y = 21.

OpenStudy (anonymous):

Answers: 24 12 36 16

OpenStudy (hoblos):

y varies directly as x means that y = kx y = 14 when x = 8 then 14 = 8k so k=14/8 = 7/4 for y=21 21 = 7/4x then x = 21*4/7 = 12